OpenStudy (anonymous):

a plot of land in the shape of a right triangle has a perimeter of 56 miles. If the length of the hypotenuse is 25 miles, find the length of two sides.

OpenStudy (anonymous):

the two legs together are \(56-25=31\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so if one leg is say \(x\) then the other one must be \(31-x\) and by pythagoras \[x^2+(31-x)^2=25^2\]

OpenStudy (anonymous):

btw we could probably guess this answer in any case since it is cooked up to give whole number solutions